OpenStudy (anonymous):

Monday night there were 33 kids and 20 adults at the movie theater. That night the box office collected $464 in ticket sales. On tuesday there were 17 kids and 10 adults for a total of $236 in ticket sales. What is the cost for 1 kid's ticket?

OpenStudy (anonymous):

it would be $8

OpenStudy (anonymous):

how do you get that?

OpenStudy (anonymous):

@Caitlynn_Faucher Remember not to just give answers as people come here to see how people get the answer. If they just wanted the answer, they would look in the back of the book.

OpenStudy (anonymous):

x=cost for 1 kid's ticket y=cost for 1 adult's ticket 33x+20y = 464 17x+10y = 236 ]*2 33x+20y = 464 34x+20y = 472 subtract the 2 equations -x = -8 so x=$8

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

Welcome!